Circular No. 9123

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 DG_{56} (WISE) D. J. Tholen, University of Hawaii (UH), reports that three 300-s ex-

posures of 2010 DG_{56}, an object with a retrograde orbit discovered by the WISE satellite (discovery observation tabulated below; cf. MPEC 2010-D79, MPS 314003-314004), taken by M. Micheli, J. Gazak, and Tholen with the UH 2.24-m reflector on Mar. 5.3 UT, show a very faint tail extending at least 3′′ to the east-southeast, with the coma being ∼ 0′′.3 larger than stars of similar brightness; Tholen adds that earlier Mauna Kea images from Feb. 20 and 21 show no obvious asymmetry. Following Tholen’s report, A. Mainzer (Jet Propulsion Laboratory) notes that WISE 12-μm images show the object to have a “soft” condensed nuclear region of size ∼ 5′′–6′′ with a ‘hazy’ coma extending ∼ 10′′ around the central core.

2010 UT α_{2000} δ_{2000} Observer
Feb. 18.00105 4^{h}30^{m}29.^{s}32 −2^{o}42′04′′.7 WISE

The recent astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2010-E43.

T = 2010 May 13.5729 TT  ω = 317.^{o}0731
Ω = 4.0718 }2000.0
q = 1.612228 AU  i = 160.3957

2009 QY_{6} AND 2010 CG_{55}

More apparently asteroidal objects with retrograde orbits have been found recently, including 2009 QY_{6} (T = 2009 Oct. 4, q = 2.07 AU, i = 138^{o}, P = 44 years; cf. MPO 167745); and 2010 CG_{55} (T = 2010 Oct. 8, q = 2.86 AU, i = 146^{o}, P = 93 years; cf. MPO 171878). There are now 27 such objects known (all but two having perihelion distances inside the orbit of Jupiter, and 22 having q < 4 AU), including numbered objects (20461) = 1999 LD_{31} (cf. IAUC 7208) and (65407) 2002 RP_{120}.

COMET 29P/SCHWASSMANN-WACHMANN Additional visual total-magnitude estimates (cf. IAUC 9116): Feb.

10.50, 10.7 (C. Wyatt, Walcha, NSW, Australia, 25-cm reflector); 19.61, 10.6 (M. Mattiazzo, Castlemaine, Victoria, Australia, 20-cm reflector); Mar. 1.81, 10.8 (J. J. Gonzalez, Asturias, Spain, 20-cm reflector); 5.88, 10.6 (U. Pilz, Leipzig, Germany, 32-cm reflector); 8.89, 11.2 (Gonzalez).

2010 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green